The perimeter P of a rectangle with length l and width w is P= 2L + 2W. What equation gives the length of a rectangle in terms of the perimeter and width? What is the length of a rectangle that has a perimeter of 25.4 meters and a width of 4.2 meters?
Answer by mananth(16946)

p=2L+2W
p-2W=2L
(p-2W)/2= L
(25.4-2*4.2)/2=L
7/2=L
3.5m = L